OpenStudy (anonymous):

Please help. Medal/Fan/Testomonial Of the following, which is the solution to 2x2 + 3x = –5? the quantity negative three plus or minus i times the square root of thirty one over four the quantity negative three plus or minus i times the square root of thirty one over two the quantity three plus or minus i times the square root of thirty one over four the quantity negative three plus or minus i times the square root of fifty over four

OpenStudy (anonymous):

\[-3 \pm i \sqrt{31} all \over 4\] is the first option

OpenStudy (amistre64):

have you considered the quadratic formula?

OpenStudy (anonymous):

whats that?

OpenStudy (anonymous):

I know the distance formula

OpenStudy (amistre64):

its the formulaic way to determine the solutions to a quadratic equations - its the shortcut of completing the square

OpenStudy (anonymous):

can you help me? Im really stuck

OpenStudy (amistre64):

what methods does your material cover?

OpenStudy (anonymous):

distance formula

OpenStudy (anonymous):

I know that the answer is not B

OpenStudy (amistre64):

as far as i can recollect, the distance formula is not useful in determining the solutions to a quadratic equation.

OpenStudy (anonymous):

then what do I do? thats all that they taught

OpenStudy (amistre64):

given a quadratic equation of the form:\[ax^2 + bx + c = 0\] we can utilize the quadratic formula:\[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (amistre64):

this formula should have been in your course work, or the method known as completing the square.

OpenStudy (anonymous):

oh, the quadratic formula is the distance formula I was talking about. Sorry, I got them confused

OpenStudy (amistre64):

distance formula: \[d=\sqrt{x^2 + y^2}\] yeah, not quite the same thing :)

OpenStudy (anonymous):

oops, well then thats my fault. Im sorry. Can you help me with the quadratic formula since that is what I meant all along?

OpenStudy (amistre64):

well, can you rewrite your problem in: ax^2 + bx + c = 0 format?

OpenStudy (anonymous):

2x^2 + 3x +5 = 0

OpenStudy (anonymous):

where 2x^2 = a b = 3x and c= 5

OpenStudy (amistre64):

a,b,c are the coefficients, not the terms themselves. [2]x^2 + [3]x +[5] = 0 a b c

OpenStudy (anonymous):

right...okay, so what now?

OpenStudy (amistre64):

plug and play of course ... everywhere in the formula that you see an a at .. replace it with a 2 everywhere in the formula that you see an b at .. replace it with a 3 everywhere in the formula that you see an c at .. replace it with a 5

OpenStudy (anonymous):

\[-3 \pm \sqrt 3^2 - 4(2)(5) \over 4\]

OpenStudy (anonymous):

which is \[-3 \pm \sqrt -9 - 40 \over 4\]

OpenStudy (anonymous):

I mean regular 9 not -9

OpenStudy (amistre64):

we know its not B (assuming you tried that first and it said it was wrong) and we need a -3 starting off, that narrows it to A or D so the real question is: is the difference between 9 and 40 bigger or smaller than 40?

OpenStudy (anonymous):

smaller

OpenStudy (anonymous):

so the answer is A?

OpenStudy (amistre64):

A is what we get to yes \[\frac{-3\pm\sqrt{9-40}}{4}\]

OpenStudy (anonymous):

Thank you so much! You were a great help! I gave you a medal and a fan

OpenStudy (amistre64):

youre welcome

OpenStudy (usukidoll):

correct. since 9-40 = -31 and we have a negative inside the radical that negative becomes i so we have \[\frac{-3\pm\sqrt{31}i}{4}\]

OpenStudy (anonymous):

thank you too UsukiDoll...

OpenStudy (usukidoll):

:) you're welcome.

OpenStudy (anonymous):

I have another small question

OpenStudy (anonymous):

If something has a vertex of (-4,-1) would the formula be (x + 4)2 − 1

OpenStudy (amistre64):

almost

OpenStudy (anonymous):

what if that is all the information I have. If they only give me the vertex is that correct? I know that there is supposed to be an 'a' in front of the parenthesis but that is not an option

OpenStudy (amistre64):

(x+4)^2 - 1 does have a vertex at (-4,-1) but so does: -3(x+4)^2 - 1 have a vertex at (-4,-1) and: 12(x+4)^2 - 1 have a vertex at (-4,-1)

OpenStudy (amistre64):

y = a(x+4)^2 - 1 have a vertex at (-4,-1) for any real value of 'a'

OpenStudy (amistre64):

except maybe a=0 ... but eh

OpenStudy (anonymous):

the first question was right and the one we did just now is right too

OpenStudy (anonymous):

thank you both so much!

OpenStudy (amistre64):

good luck